A standard deck of 52 cards is shuffled. What is the probability that
(a) the top card is a heart?
(b) the top card is a 5?
(c) the top card is the King of diamonds?
(d) the top card is a face card?
(e) the top card red and the second card is black?
(f) the top card is a 3 and the second card is an 8?
(g) the top two cards are Aces?
(h) the top three cards are Spades?

Answer by pie90010(44)

A. 13/52 = .25 = 25% chance
B. 4/52 = .076 = about A 7.5% chance
C. 1/52 = .019 = about 1.9% chance
D. 12/52 = .23 = 23% chance
E. (26/52)(26/51) = .549 = about a 54.9% chance
F. (4/52)(4/51) = .0060 = .06% chance
G. (4/52)(3/51) = .0045 = .04% chance
H. (13/52)(12/51)(11/50) = .0129 = about 1.2% chance